Summary

It is very cloudy tonight, we lost 5 hours due to the bad weather, 2.5 hours due to the telescope, which is the pointing, we can not find the guide star for the first two fields, and 1.0 hours due to the instrument, which it can not put the mask in and out of the beam. So we open 6.0 hours and got 50 minutes scientific data. macs0451 J H K image 30 min J095858 H+K spectral 20 min

12:30 Sunset, Cloudy, we do not open.

03:37 We are open. There is still some thin cloud. Pointing ~10" moving in CA direction Collimation: try BS9108 can not find guide star. Go to BS9107, then we can find the guide star, and the star is right in the centre. The collimation takes a little while, as the seeing is not good 1"

3:58 Go to the first science field: SN2008hy. Condition: 1" seeing, thin cloud Again we can not find the guide star, and we redo pointing: ~50" correction in CA direction Go back to the science field, still can not find guide star, take a short exposure, can see some stars Go back to check the pointing, only ~2" correction. Go to the science field, still can not. Go to the persson star, I can not find the guide star. Check pointing 20" in CA direction. The pointing is very bad, making us can not find the guide star.

*05:45 * Go to macs0744, the guide star is at the edge of the guider. Go to macs0451, do pointing, go back to macs0451 J imaging # 0011-#0020 Seeing is 0.8, thin cloud H band images #0021-#0030 Ks band images #0031-#0040, the seeing is 1.0" and thin cloud The last few Ks images are bad, because the could is getting thick.

06:57 We are closed due to the cloud

10:43 We are open, some thin clouds in the sky 1.2" seeing Go to J095858 H+K spectra, the first one is #50 We can see the target. Lamp+flat

11:43 Move to the next target cosmos2, can not find guide star, Check pointing, WT10_220 is a good pointing star for cosmos field. Mask can not be put to turnout, and there is no error message. And it is getting cloudy.

12:09 Telescope is closed.

12:14 The instrument team move the mask from Germany. Then, we still can not switch the mask We restart the instrument control GUI, then it is happy. And then we meet an error, peter is working on it

12:27 It is clear.

12:38 We are open, go to cosmos2, do pointing.

12:44 We met the some error

-- FuyanBian - 08 Jan 2010
Topic revision: r6 - 08 Jan 2010, FuyanBian
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback